OpenStudy (anonymous):

WRITE AN EQUATION OF A LINE IN STANDARD FORM WITH THE GIVEN SLOPE THROUGH THE GIVEN POINT -2/3(6,9)

OpenStudy (anonymous):

The slope is the change in y over the change in x. \[(y _{1}-y _{2})/(x _{1}-x_{2})\] you know two of the x and y values already. \[(9-y _{2})/(6-x_{2})\] You also know the slope; -2/3 So you know \[(9-y )/(6-x)=-2/3\] Then rewrite as an equation. You can do it.

OpenStudy (anonymous):

use general form y-y1 = m(x-x1) y-9 = -2/3 (x - 6) y-9 = -2/3x + 4 y + 2/3x = 13 3y + 2x = 39